OpenStudy (anonymous):

Use a arithmetic sequence, arithmetic series, geometric sequence, or geometric series to solve? DO not need to solve just which formula should I use? Thanks! Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August?

OpenStudy (anonymous):

Definately geometric series

OpenStudy (anonymous):

I said geometric sequence.

OpenStudy (anonymous):

I used this forula sn= a1-a1rn/1-r

OpenStudy (anonymous):

Summation of 5(2)^n

OpenStudy (amistre64):

An = 2*A{n-1}

OpenStudy (anonymous):

actually question just asks how much was deposited in august, not total. so it is \[5\times 2^6\] if i counted months correctly

OpenStudy (anonymous):

oh yeah, then sequence if you don't want the accumulation

OpenStudy (anonymous):

I came up with 320, but I don't need to submit this answer just the formula.

OpenStudy (amistre64):

An = 2.2.A{n-2} = 2.2.2.A{n-3} = 2^r . A{n-r} when r = n; we get to A{0} An = 2^{n}. A{0} or did i mess it up somewhere?

OpenStudy (amistre64):

how much was the deposit in august or how much in total has been depostited?

OpenStudy (anonymous):

Are we in agreement that it is a geometric sequence since I only need to know August and not the total sum?

OpenStudy (amistre64):

geo if we only want to know the amount for augusts deposit

OpenStudy (anonymous):

you lost me it is feb mar apr may june july aug 5, 10, 20, 40, 80, 160, 320

OpenStudy (anonymous):

yes geometric \[5, 5\times 2, 5\times 2^2, 5\times 2^3, ...\]

OpenStudy (amistre64):

m: f m a m j j a n: 0 1 2 3 4 5 6 A: 5 10 20 40 80 160 320

OpenStudy (anonymous):

geometric sequence

OpenStudy (anonymous):

Thanks. Great conversation!

OpenStudy (amistre64):

i tend to start at a0 instead of a1 :)